a planet of mass M moves in circular orbit of radius r0 in the gravitational potential vr=-k/r then what is the angular momentum of this planet.
1
Expert's answer
2020-09-06T17:17:03-0400

Given potential is "V = -\frac{k}{r}"

Let L be angular momentum then total effective potential will be

"V_0 = V + \frac{L^2}{2Mr^2} = -\frac{k}{r} + \frac{L^2}{2Mr^2}"


Differentiate with respect to r and differentiation will be zero at "r = r_0"


Hence, "\frac{dV}{dr} = \frac{k}{r^2} - \frac{L^2}{Mr^3}"


"\frac{dV}{dr}|_{r=r_0} = \frac{k}{r_0^2} - \frac{L^2}{Mr_0^3} = 0"


Solving it we get,

"L = \sqrt{kMr_0}"